Two satellites are in circular orbits around the earth. The orbit for satellite A is at a height of 502 km above the earth’s surface, while that for satellite B is at a height of 747 km. Find the orbital speed for (a) satellite A and (b) satellite B.
Answer #1

R = radius of earth = 6371 km

mass of earth M = 5.98*10^24 kg

hA = 502 km , hB = 747 km

orbital speed v = (GM/(R+h))^0.5

(a) vA = (6.67*10^-11*5.98*10^24/(6371000+502000))^0.5

vA = 7.618 km/s

(b) vB = (6.67*10^-11*5.98*10^24/(6371000+747000))^0.5

vB = 7.486 km/s